# QCircuitNet: A Large-Scale Hierarchical Dataset for Quantum Algorithm Design

### v1

@@ Appendix B Additional Preliminaries for Quantum Computing and Quantum Information
A quantum algorithm is composed of a series of quantum gates. By default, a quantum algorithm starts from the all-0 state $|0^{n}\rangle$. A quantum algorithm can be illustrated by its quantum gate diagram, drawn from left to right. The initial all-0 state is placed at the left side of the diagram. After that, whenever we apply a quantum gate, it is placed on the corresponding qubits, from left to right. At the end of the quantum gates, we need to measure and read the outputs, and these measurements are placed at the right side of the diagram. See Figure 4 for the quantum gate diagram of Simon’s algorithm (Simon, 1997).
[…]
Figure 4: Quantum gate diagram of Simon’s algorithm.
Superdense Coding.
Superdense coding (Bennett & Wiesner, 1992) is a quantum communication protocol that allows Alice to transmit two classical bits of information to Bob by sending only one qubit, given that they share a pair of entangled qubits. The protocol can be divided into five steps:
  1. 1.
    Preparation: Charlie prepares a maximally entangled Bell state, such as $|\beta_{00}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$.
  2. 2.
    Sharing: Charlie sends the qubit 1 to Alice and the qubit 2 to Bob. Alice and Bob can be separated by an arbitrary distance.
  3. 3.
    Encoding: Depending on the two classical bits $zx\in\{00,01,10,11\}$ that Alice wants to send, she applies the corresponding quantum gate operation to her qubit, transforming the Bell state $|\beta_{00}\rangle$ into one of the four Bell states:
[…]
    Alice achieves these transformations by applying the operation $Z^{z}X^{x}$ to her qubit, where $Z$ is the phase-flip gate, $X$ is the bit-flip gate. Specifically:
    • If $zx=00$, Alice applies $Z^{0}X^{0}=I$ (identity gate).
    • If $zx=01$, Alice applies $Z^{0}X^{1}=X$ (bit-flip gate).
    • If $zx=10$, Alice applies $Z^{1}X^{0}=Z$ (phase-flip gate).
    • If $zx=11$, Alice applies $Z^{1}X^{1}=ZX=iY$ gate.
  4. 4.
    Sending: Alice sends her qubit to Bob through a quantum channel.
  5. 5.
    Decoding: Bob applies a CNOT gate followed by a Hadamard gate to the two qubits, transforming the entangled state into the corresponding computational basis state $|zx\rangle$. By measuring the qubits, Bob obtains the two classical bits $zx$ sent by Alice.
Superdense coding exploits the properties of quantum entanglement to transmit two classical bits of information using only one qubit. The quantum circuit diagram for superdense coding is shown in Figure 5.
[…]
Figure 5: Quantum circuit diagram for superdense coding.
Quantum Teleportation.
Quantum teleportation (Bennett et al., 1993) is a technique for transferring quantum information from a sender (Alice) to a receiver (Bob) using shared entanglement and classical communication. The protocol can be described as follows:
  1. 1.
    Preparation: Telamon prepares a maximally entangled Bell state, such as $|\beta_{00}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$.
  2. 2.
    Sharing: Alice has qubit 1 in the state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, which she wants to teleport to Bob. Telamon shares qubit 2 with Alice and qubit 3 with Bob, creating the shared entangled state $|\beta_{00}\rangle_{23}$.
  3. 3.
    Encoding: Alice wants to teleport an unknown quantum state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ to Bob. She applies a CNOT gate to qubits 1 and 2, with qubit 1 as the control and qubit 2 as the target. Then, she applies a Hadamard gate to qubit 1. The resulting state of the three-qubit system is:
    $|\Psi\rangle$ $=\frac{1}{2}[|\beta_{00}\rangle(\alpha|0\rangle+\beta|1\rangle)+|\beta_{01}\rangle(\alpha|1\rangle+\beta|0\rangle)$
    $+|\beta_{10}\rangle(\alpha|0\rangle-\beta|1\rangle)+|\beta_{11}\rangle(\alpha|1\rangle-\beta|0\rangle)].$
  4. 4.
    Measurement: Alice measures qubits 1 and 2 in the Bell basis and obtains one of four possible outcomes: $|\beta_{00}\rangle$, $|\beta_{01}\rangle$, $|\beta_{10}\rangle$, or $|\beta_{11}\rangle$. This measurement collapses the three-qubit state into one of the following:
    $|\beta_{00}\rangle$ $\otimes(\alpha|0\rangle+\beta|1\rangle)$
    $|\beta_{01}\rangle$ $\otimes(\alpha|1\rangle+\beta|0\rangle)$
    $|\beta_{10}\rangle$ $\otimes(\alpha|0\rangle-\beta|1\rangle)$
    $|\beta_{11}\rangle$ $\otimes(\alpha|1\rangle-\beta|0\rangle)$
  5. 5.
    Classical Communication: Alice sends the result of her measurement (two classical bits) to Bob via a classical channel.
  6. 6.
    Reconstruction: Depending on the classical information received from Alice, Bob applies the operation $Z^{z}X^{x}$ to qubit 3, where $z$ and $x$ correspond to the two classical bits sent by Alice:
    • If Alice measured $|\beta_{00}\rangle$, she sends $zx=00$, and Bob applies $Z^{0}X^{0}=I$ (identity operation).
    • If Alice measured $|\beta_{01}\rangle$, she sends $zx=01$, and Bob applies $Z^{0}X^{1}=X$ (bit-flip).
    • If Alice measured $|\beta_{10}\rangle$, she sends $zx=10$, and Bob applies $Z^{1}X^{0}=Z$ (phase-flip).
    • If Alice measured $|\beta_{11}\rangle$, she sends $zx=11$, and Bob applies $Z^{1}X^{1}=ZX=iY$ (bit-flip and phase-flip).
    After applying the appropriate operation, Bob’s qubit 3 will be in the state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, which is the original state that Alice wanted to teleport.
The quantum circuit diagram for quantum teleportation is shown in Figure 6.
[…]
Figure 6: Quantum circuit diagram for quantum teleportation
Quantum Key Distribution.
Quantum key distribution (QKD) (Bennett & Brassard, 1984) is a secure communication protocol that allows two parties, Alice and Bob, to produce a shared random secret key, which can then be used to encrypt and decrypt messages. The security of QKD is based on the fundamental principles of quantum mechanics that measuring a qubit can change its state. One of the most well-known QKD protocols is the BB84 protocol, which works as follows:
  1. 1.
    Alice randomly generates a bit string and chooses a random basis (X or Z) for each bit. She then encodes the bits into qubits using the chosen bases and sends them to Bob through a quantum channel.
  2. 2.
    Bob measures the received qubits in randomly chosen bases (X or Z) and records the results.
  3. 3.
    Alice and Bob communicate over a public classical channel to compare their basis choices. They keep only the bits for which their basis choices coincide and discard the rest.
  4. 4.
    Alice and Bob randomly select a subset of the remaining bits and compare their values. If the error rate is below a certain threshold, they conclude that no eavesdropping has occurred, and the remaining bits can be used as a secret key. If the error rate is too high, they abort the protocol, as it indicates the presence of an eavesdropper (Eve).
The security of the BB84 protocol relies on the fact that any attempt by Eve to measure the qubits during transmission will introduce detectable errors, alerting Alice and Bob to the presence of an eavesdropper.
Appendix C Additional Experiment Results
In this section, we include detailed analysis of the experiments and additional experiment results. In Section C.1, we introduce the metrics: BLEU score, verification score, and byte perplexity, and provide a detailed analysis for the experiments on BLEU and verification score. In Section C.2, we include additional experiments on perplexity score and temperatures. In Section C.3, we present concrete cases of typical patterns observed in model outputs. In Section C.4, we analyze the data contamination phenomenon revealed by our benchmark.
C.1 Metrics
BLEU Score.
Bilingual Evaluation Understudy (BLEU) score is a metric used to evaluate the quality of machine-translated text compared to human-translated text. It measures how close the machine translation is to one or more reference translations. The BLEU score evaluates the quality of text generated by comparing it with one or more reference texts. It does this by calculating the n-gram precision, which means it looks at the overlap of n-grams (contiguous sequences of n words) between the generated text and the reference text. Originally the BLEU score ranges from 0 to 1, where 1 indicates a perfect match with the reference translations. Here rescaling the score makes it ranges from 0 to 100.
The BLEU score, originally designed for machine translation, can also be effectively used for evaluating algorithm generation tasks. Just as BLEU measures the similarity between machine-translated text and human reference translations, it can measure the similarity between a generated algorithm and a gold-standard algorithm. This involves comparing sequences of tokens to assess how closely the generated output matches the reference solution. In the context of algorithm generation, n-grams can represent sequences of tokens or operations in the code. BLEU score captures the precision of these n-grams, ensuring that the generated code aligns closely with the expected sequences found in the reference implementation.
The formula for BLEU score is given by:
$\text{BLEU}=BP\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right).$
where $BP$ is the acronym for brevity penalty, $w_{n}$ is the weight for the n-gram precision (typically $\frac{1}{N}$ for uniform weights), $p_{n}$ is the precision for n-grams. BP is calculated as:
$BP=\begin{cases}1&\text{if }c>r\\ e^{1-\frac{r}{c}}&\text{if }c\leq r\end{cases}.$
where $c$ is the length of the generated text and $r$ is the length of the reference text. Furthermore, n-gram precision $p_{n}$ is calculated as:
$p_{n}=\frac{\sum_{C\in\text{Candidates}}\sum_{n-\text{gram}\in C}\min(\text{Count}(n-\text{gram in candidate}),\text{Count}(n-\text{gram in references}))}{\sum_{C\in\text{Candidates}}\sum_{n-\text{gram}\in C}\text{Count}(n-\text{gram in candidate})}.$
This formulation ensures that the BLEU score takes into account both the precision of the generated n-grams and the overall length of the translation, providing a balanced evaluation metric.
In our experiments, the BLEU scores for various quantum algorithm design tasks are illustrated in Figure 3. This figure not only displays the average performance of each model but also highlights the differences in performance across individual quantum algorithm tasks. The first notable observation is that the figure clearly demonstrates the varying levels of difficulty among quantum algorithms. For example, models achieve higher BLEU scores on tasks such as Bernstein-Vazirani and Deutsch-Jozsa, whereas they perform significantly worse on tasks like Grover, phase estimation, and quantum Fourier transform. This indicates that the former tasks are considerably easier than the latter ones. Another significant observation is that most models score higher in a five-shot prompt compared to a one-shot prompt, which confirms the large language models’ ability to improve performance through contextual learning.
Similar patterns are observed in oracle construction tasks, as illustrated in Figure 3. The figure highlights that the Diffusion Operator task is notably more challenging than the Grover oracle construction task. Interestingly, we found that adding more in-context examples actually reduced the performance of the Phi-3-medium-128k-instruct and Mistral-7B-v0.3 models. This decline in performance could be attributed to the significant differences between each oracle construction task, which may be too out-of-distribution. Consequently, the additional examples might cause the models to overfit to the specific examples provided in the context, rather than generalizing well across different tasks.
Detailed Analysis of Verification Score.
In addition to evaluating the BLEU score, we conducted an experiment to measure the correctness of the machine-generated algorithms, and the results are shown in Table 1. By running a verification function, we discovered that phase estimation and the swap test are significantly more challenging than other problems, leading most models to score -1 (indicating they cannot even generate the correct syntax). Notably, the BLEU score for the swap test is above average compared to other algorithms, yet almost none of the models produced a correct algorithm. This discrepancy highlights a critical limitation of using BLEU as a metric for algorithm evaluation. BLEU measures average similarity, but even a single mistake in an algorithm can render it entirely incorrect, thus failing to capture the true accuracy and functionality of the generated algorithms. Another important finding is that in a five-shot setting, GPT-4 and GPT-3.5 surpass all other models by a large margin. This demonstrates their exceptional capabilities, particularly in long-context comprehension and in-context learning. These models not only excel in understanding and generating text based on minimal examples but also maintain high performance over extended sequences, highlighting their advanced architecture and training methodologies.
The verification results of the oracle construction task, as shown in Table 2, confirm our previous conclusions. In the five-shot setting, GPT-4 and GPT-3.5 consistently outperform all other models. Additionally, this table highlights the inconsistency between BLEU scores and verification scores. For instance, while the Diffusion Operator task achieves the lowest BLEU score, it is the Grover oracle construction that receives the lowest verification score. This discrepancy suggests that BLEU scores may not fully capture the performance of models in certain complex tasks, and it is necessary to include verification score as a comprehensive evaluation.
Byte Perplexity.
Perplexity is a measure of how well a probability distribution or a probabilistic model predicts a sample. In the context of language models, it quantifies the uncertainty of the model when it comes to predicting the next element in a sequence. Byte perplexity specifically deals with sequences of bytes, which are the raw binary data units used in computer systems. For our purposes, we consider byte perplexity under UTF-8 encoding, a widely used character encoding standard that represents each character as one or more bytes.
For a given language model, let $p(x_{i}|x_{<i})$ be the probability of the $i$-th byte $x_{i}$ given the preceding bytes $x_{<i}$. If we have a sequence of bytes $x=(x_{1},x_{2},\ldots,x_{N})$, the perplexity $\textrm{PPL}(x)$ of the model on this sequence is defined as:
$\textrm{PPL}(x)=2^{-\frac{1}{N}\sum_{i=1}^{N}\log_{2}p(x_{i}|x_{<i})}.$
A notable feature of byte perplexity is that, it does not rely on any specific tokenizer, making it versatile for comparing different models. Therefore, byte perplexity can be used to measure the performance in quantum algorithm generation tasks. In such tasks, a lower byte perplexity indicates a better-performing model, as it means the model is more confident in its predictions of the next byte in the sequence.
C.2 Additional Experiments on PPL Score and Temperatures
The Byte Perplexity results, shown in Figure 7, provide valuable insights into the performance of our model. Evaluated in a zero-shot setting, byte perplexity trends closely mirror those observed with BLEU scores. This alignment suggests that our model’s predictive capabilities are consistent across Perplexity and BLEU evaluation metrics. Specifically, in the context of quantum algorithm design tasks, the results indicate that the Bernstein-Vazirani and Deutsch-Jozsa algorithms are relatively straightforward for the model, whereas the Simon algorithm presents greater difficulty. This differentiation highlights the varying levels of complexity inherent in these quantum algorithms.
Figure 7: Benchmarking algorithm design and oracle construction in perplexity scores.
Regarding the counter-intuitive phenomenon where the performance on Clifford and universal random circuits decreases after fine-tuning, we conducted additional experiments and fine-tuned the model on 4,800 samples specifically for the Clifford task. Upon closer inspection, we observed that the model more frequently generated outputs with infinite loops and increased monotony, often producing repetitive gate patterns and repeatedly cycling over the same qubit after fine-tuning. We further conducted experiments with different "temperature" parameters, which control the randomness of predictions. Formally, let $T>0$ be the temperature, $z_{i}$ be the raw score for token $i$, the probability for token $i$ is computed as $p_{i}=\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}$. Typically, lower temperatures make the model more conservative, while higher temperatures flatten the distribution, increasing the likelihood of generating originally less probable sequences. The results are shown in Table 4:
Table 4: Clifford Model Fine-Tuning Results Across Different Temperature Settings
Model Setting Temperature BLEU Verification
Llama3 few-shot(5) 0 13.3796(±0.9508) -0.6582(±0.0360)
0.2 12.5688(±0.8276) -0.6526(±0.0372)
1 53.0431(±3.8422) -0.1914(±0.0361)
Llama finetune 0 7.6261(±0.3433) -0.8895(±0.0247)
0.2 13.8714(±0.6536) -0.7873(±0.0306)
1 32.5241(±2.0548) -0.2072(±0.0358)
One possible explanation for this counter-intuitive result lies in the challenge of encoding quantum state vectors within a language model. In the problem description, the target quantum state is represented by a complex vector with four decimal places of precision, where the dimension scales as with the number of qubits . It is a well-known fact that LLMs generally struggle with very long floating-point numbers, which might contribute to the observed performance decline.
Another potential reason could be overfitting during fine-tuning, particularly for tasks that require high output diversity. The varying degrees of intrinsic difficulty and the amount of relevant pre-training knowledge across different tasks likely played a role. Oracle constructions are relatively simple for the model to learn. For example, in the Bernstein-Vazirani algorithm, the model only needs to apply a CNOT gate at positions corresponding to ’1’ bits. In contrast, the random circuits in the Clifford and Universal tasks involve more general and complex quantum state transformations, making them significantly more challenging. These tasks are also less common during pre-training, which could have hindered the model’s ability to generalize without overfitting. This challenge is one of the reasons we initially considered a few-shot learning approach to be suitable.
While these are plausible hypotheses, we acknowledge that further investigation is required to draw definitive conclusions. We consider this an intriguing topic that warrants additional research.
C.3 Case Study
After carefully examining the model’s output, we observed several interesting patterns. We present a series of case studies to illustrate these observations and provide possible explanations.
Low Score for GPT-4o in One-Shot Setting.
At first glance, it is surprising that GPT-4o performs poorly on many quantum algorithms in the algorithm design task in the one-shot setting compared to Llama3-8B. Given that Llama3-8B has a relatively smaller parameter scale, the results should have been the other way around. A closer examination of the model’s output reveals the potential reason: while Llama3-8B closely mimics the input examples, GPT-4o tends to improvise, resulting in outputs that are not well captured by the current syntax support. Here are several concrete examples.
This is the OpenQASM 3.0 code output for the W state with $n=7$. In this code, GPT-4o uses the advanced "for" loop syntax newly introduced in OpenQASM 3.0 to create the circuit. Although the code fails to produce the W state, it is syntactically correct. However, the Qiskit.qasm3 import module, which converts OpenQASM 3.0 files to QuantumCircuit objects and is used in our verification function to check the correctness of the syntax of output OpenQASM codes, is still in the experimental stage and does not support many of OpenQASM 3.0’s advanced features. As a result, GPT-4o’s use of these features causes the code to fail syntax validation, resulting in a score of -1.
OPENQASM 3.0;
include "stdgates.inc";
qubit[7] q;
h q[0];
for i in[1:6] {
cx q[i-1], q[i];
Listing 8: OpenQASM 3.0 Code output by GPT-4o for W state with $n=7$.
@@ C.3 Case Study
In contrast to its relatively low score in the one-shot setting, GPT-4o achieves the best overall performance in the five-shot setting for both oracle construction and algorithm design tasks.
Unlike in the one-shot setting, where GPT-4o frequently uses advanced OpenQASM 3.0 features, it learns from the examples in context and uses simpler syntax in the five-shot setting. For the W state task, here is an output in the five-shot setting:
OPENQASM 3.0;
include "stdgates.inc";
qubit[3] q;
h q[0];
cx q[0], q[1];
h q[1];
cx q[1], q[2];
h q[2];
Listing 10: OpenQASM 3.0 Code output by GPT-4o for W state with $n=3$.
This output avoids the use of "for" loops and successfully passes the syntax validation test, although it still cannot generate the W state correctly.
In addition to adapting to plain syntax through in-context learning, GPT-4o achieves outstanding performance on more complicated tasks such as phase estimation. Here is the model output of GPT-4o on the phase estimation task with qubit number $n=2$.
OPENQASM 3.0;
include "stdgates.inc";
include "oracle.inc";
bit[2] c;
qubit[2] q;
Psi q[0];
h q[1];
CU_0 q[0], q[1];
h q[1];
c[0] = measure q[1];
c[1] = measure q[0];
Listing 11: OpenQASM 3.0 Code output by GPT-4o for Phase Estimation with $n=2$.
from Qiskit import transpile
def run_and_analyze(circuit, aer_sim):
"""Run the circuit and analyze the result."""
circ = transpile(circuit, aer_sim)
result = aer_sim.run(circ, shots=1).result()
counts = result.get_counts()
phase_str = list(counts.keys())[0]
phase = int(phase_str, 2) / 2**2
return phase
Listing 12: Post-processing code output by GPT-4o for Phase Estimation with $n=2$.
This suite of OpenQASM 3.0 circuits and post-processing functions successfully outputs the phase within the required precision for the test case, resulting in an impressive verification score of 1.0. Despite the small number of qubits and differences from the reference implementation, the accuracy achieved is noteworthy.
These phenomena reflect that GPT-4o has impressive in-context learning abilities and overall better capabilities in designing and implementing quantum algorithms.
C.4 Data Contamination
In our experiments, we observed an interesting phenomenon of performance discrepancy between implementing quantum algorithms with Qiskit codes for the general setting and with OpenQASM codes for a concrete case.
We provided the following prompt:
Prompt for Bernstein-Vazirani Problem with Qiskit Please write the code with Qiskit implementing Bernstein-Vazirani Algorithm with qubit number n = 5.
The model answered with the following text and codes:
Text Answer by GPT4o Here’s a Python code using Qiskit to implement the Bernstein-Vazirani algorithm with $n=5$ qubits. The secret string used in this example is $s=10101$ (you can change this to any string of your choice).
from qiskit import QuantumCircuit, Aer, transpile, assemble
from qiskit.visualization import plot_histogram
# Secret string s for the Bernstein-Vazirani problem
s = ’10101’
# Number of qubits
n = len(s)
# Create a Quantum Circuit with n+1 qubits (one ancillary qubit)
bv_circuit = QuantumCircuit(n+1, n)
# Apply Hadamard gates to all qubits except the ancillary qubit
bv_circuit.h(range(n))
# Apply X and Hadamard gate to the ancillary qubit
bv_circuit.x(n)
bv_circuit.h(n)
# Construct the Oracle
for i, bit in enumerate(s):
if bit == ’1’:
bv_circuit.cx(i, n)
# Apply Hadamard gates to the first n qubits
bv_circuit.h(range(n))
# Measure the first n qubits
bv_circuit.measure(range(n), range(n))
# Visualize the circuit
bv_circuit.draw(’mpl’)
# Run the circuit using Qiskit Aer simulator
simulator = Aer.get_backend(’qasm_simulator’)
t_qc = transpile(bv_circuit, simulator)
qobj = assemble(t_qc)
result = simulator.run(qobj).result()
# Get the result counts and plot the histogram
counts = result.get_counts()
plot_histogram(counts)
Listing 13: Qiskit code for Bernstein-Vazirani Problem generated by GPT4o in the zero-shot setting.
This code, although adopts an older version of Qiskit syntax, can be compiled and run successfully with the correct answer. At first glance, this indicates that GPT4o has fully grasped the mechanism of the Bernstein-Vazirani Algorithm and the syntax of Qiskit. Similar performances can be observed with OpenQASM codes:
// OpenQASM 2.0
OPENQASM 2.0;
include "qelib1.inc";
// Declare 5 qubits for input and 1 ancilla qubit
qreg q[6];
creg c[5]; // classical register for measuring the result
// Step 1: Apply Hadamard gates to the input qubits
h q[0];
h q[1];
h q[2];
h q[3];
h q[4];
// Apply Hadamard to the ancillary qubit and set it to |1>
x q[5];
h q[5];
// Step 2: Oracle implementation (CNOT gates based on secret string s = 10110)
cx q[0], q[5]; // Secret string bit 1
// No cx for q[1] because the secret string bit is 0
cx q[2], q[5]; // Secret string bit 1
cx q[3], q[5]; // Secret string bit 1
// No cx for q[4] because the secret string bit is 0
// Step 3: Apply Hadamard gates to the input qubits again
h q[0];
h q[1];
h q[2];
h q[3];
h q[4];
// Step 4: Measure the qubits to get the result
measure q[0] -> c[0];
measure q[1] -> c[1];
measure q[2] -> c[2];
measure q[3] -> c[3];
measure q[4] -> c[4];
Listing 14: OpenQASM code for Bernstein-Vazirani Problem generated by GPT4o.
However, if we ask the GPT4o model to implement the code for a given secret string, for example $s=000101$, the model fails to apply the CX gate on the correct qubits.
OPENQASM 3.0;
include "stdgates.inc";
gate Oracle _gate_q_0, _gate_q_1, _gate_q_2, _gate_q_3, _gate_q_4, _gate_q_5, _gate_q_6 {
// Apply CX gates according to the secret string "000101"
cx _gate_q_2, _gate_q_6;
cx _gate_q_5, _gate_q_6;
bit[6] c;
qubit[7] q;
// Apply Hadamard gates to the first 6 qubits
h q[0];
h q[1];
h q[2];
h q[3];
h q[4];
h q[5];
// Prepare the ancilla qubit in the |1> state and apply a Hadamard gate
x q[6];
h q[6];
// Call the oracle
Oracle q[0], q[1], q[2], q[3], q[4], q[5], q[6];
// Apply Hadamard gates to the first 6 qubits again
h q[0];
h q[1];
h q[2];
h q[3];
h q[4];
h q[5];
// Measure the first 6 qubits
c[0] = measure q[0];
c[1] = measure q[1];
c[2] = measure q[2];
c[3] = measure q[3];
c[4] = measure q[4];
c[5] = measure q[5];
Listing 15: OpenQASM code for Bernstein-Vazirani Problem with s = 000101 in few-shot setting.
We reminded the model of the mechanism with the following prompt:
Reminder Prompt Remember that in the construction of oracles for Bernstein-Vazirani Problem, you should apply cx gates where the control qubits are where the string bit equals 1, and the target qubit is the ancillary qubit.
There was no improvement. We further asked the model about the indices of the bits in the string which equal 1. It provided the following answer:
GPT4o’s answer for counting indices In the secret string s="000101", the bits that equal 1 are found at the following indices (0-based): Index 2 Index 5 These indices correspond to the positions where the string has a ’1’.
This reveals the incapability for GPT models to count the bit of $1$ and apply corresponding gates to construct the oracle. Considering that Qiskit has implemented tutorial codes for several representative quantum algorithms, the seemingly outstanding performance of quantum algorithm code generation might be a deceptive result caused by data contamination. Therefore, only a carefully designed quantum algorithm dataset can avoid the effects of data contamination, allowing for an effective evaluation of the model’s genuine capability in quantum algorithm design and implementation. This dataset is also meaningful for testing general AI code generation and syntax learning, where no existing AI dataset could substitute us.

### v2

@@ Appendix B Preliminaries for Quantum Computing and Quantum Information
A quantum algorithm is composed of a series of quantum gates. By default, a quantum algorithm starts from the all-0 state $|0^{n}\rangle$. A quantum algorithm can be illustrated by its quantum gate diagram, drawn from left to right. The initial all-0 state is placed at the left side of the diagram. After that, whenever we apply a quantum gate, it is placed on the corresponding qubits, from left to right. At the end of the quantum gates, we need to measure and read the outputs, and these measurements are placed at the right side of the diagram. See Figure LABEL:fig:Simon for the quantum gate diagram of Simon’s algorithm [simon1997power].
@@
In contrast to its relatively low score in the one-shot setting, GPT-4o achieves the best overall performance in the five-shot setting for both oracle construction and algorithm design tasks. Unlike in the one-shot setting, where GPT-4o frequently uses advanced OpenQASM 3.0 features, it learns from the examples in context and uses simpler syntax in the five-shot setting. For the W state task, here is an output in the five-shot setting: OPENQASM 3.0; include "stdgates.inc"; qubit[3] q; h q[0]; cx q[0], q[1]; h q[1]; cx q[1], q[2]; h q[2]; Listing 10: OpenQASM 3.0 Code output by GPT-4o for W state with $n=3$. This output avoids the use of "for" loops and successfully passes the syntax validation test, although it still cannot generate the W state correctly. In addition to adapting to plain syntax through in-context learning, GPT-4o achieves outstanding performance on more complicated tasks such as phase estimation. Here is the model output of GPT-4o on the phase estimation task with qubit number $n=2$. OPENQASM 3.0; include "stdgates.inc"; include "oracle.inc"; bit[2] c; qubit[2] q; Psi q[0]; h q[1]; CU_0 q[0], q[1]; h q[1]; c[0] = measure q[1]; c[1] = measure q[0]; Listing 11: OpenQASM 3.0 Code output by GPT-4o for Phase Estimation with $n=2$. from qiskit import transpile def run_and_analyze(circuit, aer_sim): """Run the circuit and analyze the result.""" circ = transpile(circuit, aer_sim) result = aer_sim.run(circ, shots=1).result() counts = result.get_counts() phase_str = list(counts.keys())[0] phase = int(phase_str, 2) / 2**2 return phase Listing 12: Post-processing code output by GPT-4o for Phase Estimation with $n=2$. This suite of OpenQASM 3.0 circuits and post-processing functions successfully outputs the phase within the required precision for the test case, resulting in an impressive semantic score of 1.0. Despite the small number of qubits and differences from the reference implementation, the accuracy achieved is noteworthy. These phenomena reflect that GPT-4o has impressive in-context learning abilities and overall better capabilities in designing and implementing quantum algorithms.
C.3.2 Counting Error
If we ask the GPT-4o model to implement the code of Bernstein-Vazirani Problem for a given secret string, for example $s=000101$, the model fails to apply the CX gate on the correct qubits. OPENQASM 3.0; include "stdgates.inc"; gate Oracle _gate_q_0, _gate_q_1, _gate_q_2, _gate_q_3, _gate_q_4, _gate_q_5, _gate_q_6 { // Apply CX gates according to the secret string "000101" cx _gate_q_2, _gate_q_6; cx _gate_q_5, _gate_q_6; } bit[6] c; qubit[7] q; // Apply Hadamard gates to the first 6 qubits h q[0]; h q[1]; h q[2]; h q[3]; h q[4]; h q[5]; // Prepare the ancilla qubit in the |1> state and apply a Hadamard gate x q[6]; h q[6]; // Call the oracle Oracle q[0], q[1], q[2], q[3], q[4], q[5], q[6]; // Apply Hadamard gates to the first 6 qubits again h q[0]; h q[1]; h q[2]; h q[3]; h q[4]; h q[5]; // Measure the first 6 qubits c[0] = measure q[0]; c[1] = measure q[1]; c[2] = measure q[2]; c[3] = measure q[3]; c[4] = measure q[4]; c[5] = measure q[5]; Listing 13: OpenQASM code for Bernstein-Vazirani Problem with s = 000101 in few-shot setting. We reminded the model of the mechanism with the following prompt: Reminder Prompt Remember that in the construction of oracles for Bernstein-Vazirani Problem, you should apply cx gates where the control qubits are where the string bit equals 1, and the target qubit is the ancillary qubit. There was no improvement. We further asked the model about the indices of the bits in the string which equal 1. It provided the following answer: GPT-4o’s answer for counting indices In the secret string s="000101", the bits that equal 1 are found at the following indices (0-based): Index 2 Index 5 These indices correspond to the positions where the string has a ‘1’. This reveals the incapability for GPT models to count the bit of $1$ and apply corresponding gates to construct the oracle.
C.3.3 Data Contamination
In our experiments, we observed an interesting phenomenon of performance discrepancy between implementing quantum algorithms with Qiskit codes for the general setting and with OpenQASM codes for a concrete case. We provided the following prompt: Prompt for Bernstein-Vazirani Problem with Qiskit Please write the code with Qiskit implementing Bernstein-Vazirani Algorithm with qubit number n = 5. The model answered with the following text and codes: Text Answer by GPT-4o Here’s a Python code using Qiskit to implement the Bernstein-Vazirani algorithm with $n=5$ qubits. The secret string used in this example is $s=10101$ (you can change this to any string of your choice). from qiskit import QuantumCircuit, Aer, transpile, assemble from qiskit.visualization import plot_histogram # Secret string s for the Bernstein-Vazirani problem s = ’10101’ # Number of qubits n = len(s) # Create a Quantum Circuit with n+1 qubits (one ancillary qubit) bv_circuit = QuantumCircuit(n+1, n) # Apply Hadamard gates to all qubits except the ancillary qubit bv_circuit.h(range(n)) # Apply X and Hadamard gate to the ancillary qubit bv_circuit.x(n) bv_circuit.h(n) # Construct the Oracle for i, bit in enumerate(s): if bit == ’1’: bv_circuit.cx(i, n) # Apply Hadamard gates to the first n qubits bv_circuit.h(range(n)) # Measure the first n qubits bv_circuit.measure(range(n), range(n)) # Visualize the circuit bv_circuit.draw(’mpl’) # Run the circuit using Qiskit Aer simulator simulator = Aer.get_backend(’qasm_simulator’) t_qc = transpile(bv_circuit, simulator) qobj = assemble(t_qc) result = simulator.run(qobj).result() # Get the result counts and plot the histogram counts = result.get_counts() plot_histogram(counts) Listing 14: Qiskit code for Bernstein-Vazirani Problem generated by GPT-4o in the zero-shot setting. This code, although adopts an older version of Qiskit syntax, can be compiled and run successfully with the correct answer. Similar performances can be observed with OpenQASM codes: // OpenQASM 2.0 OPENQASM 2.0; include "qelib1.inc"; // Declare 5 qubits for input and 1 ancilla qubit qreg q[6]; creg c[5]; // classical register for measuring the result // Step 1: Apply Hadamard gates to the input qubits h q[0]; h q[1]; h q[2]; h q[3]; h q[4]; // Apply Hadamard to the ancillary qubit and set it to |1> x q[5]; h q[5]; // Step 2: Oracle implementation (CNOT gates based on secret string s = 10110) cx q[0], q[5]; // Secret string bit 1 // No cx for q[1] because the secret string bit is 0 cx q[2], q[5]; // Secret string bit 1 cx q[3], q[5]; // Secret string bit 1 // No cx for q[4] because the secret string bit is 0 // Step 3: Apply Hadamard gates to the input qubits again h q[0]; h q[1]; h q[2]; h q[3]; h q[4]; // Step 4: Measure the qubits to get the result measure q[0] -> c[0]; measure q[1] -> c[1]; measure q[2] -> c[2]; measure q[3] -> c[3]; measure q[4] -> c[4]; Listing 15: OpenQASM code for Bernstein-Vazirani Problem generated by GPT-4o. This phenomenon contrasts with the results in Appendix C.3.2. Considering that Qiskit has implemented tutorial codes for several representative quantum algorithms, this seemingly outstanding performance might be a deceptive result caused by data contamination. Therefore, only a carefully designed quantum algorithm dataset can avoid the effects of data contamination, allowing for an effective evaluation of the model’s genuine capability in quantum algorithm design and implementation. This dataset is also meaningful for testing general AI code generation and syntax learning, where no existing AI dataset could substitute us.